News

Network status: maintenance at CPT; Spectral camera image reprocessing

Here's an update on recent LCOGT activities.

The air conditioning unit in the front server room at LCOGT's HQ is awaiting repair. Our IT team has done an excellent job of keeping as-many-servers-as-possible running, and we hope that you haven't noticed any negative effects.
 
With the help of users, we discovered a problem with the bias subtraction in images from the Spectral camera (fs02) on the 2-meter telescope on Haleakala. The problem affected images between October 6 and November 22. We are reprocessing those images and replacing the images in the archive.
 
Two members of our Telescope Operations team are currently at the Sutherland site (CPT). They are there primarily to make infrastructure improvements in preparation for the installation of Sinistro cameras, but they will be doing other maintenance work as well. Their work is not expected to affect nightly science observations, but if observations must be suspended, we'll let you know.
 
Nikolaus
  
Share on Twitter Share on Facebook